Reports 1-1 of 1 Clear search Modify search
AdV-DET (External bench)
goauty, masserot, was - 15:21 Thursday 03 April 2025 (66498) Print this report
Preparatory works for EDB_OMC1 installation

Update of the Acl tasks order on the SDB_EDB_rtpc(rtpc1)

Today on the SDB_EDB_rtpc, the task repartition on the 3 RT_CPUs is shown in the following table

 RT_CPU2 RT_CPU1  RT_CPU0
 SDB1_OMC - task id 0  SDB1_Quadrants - task id 4  SDB2_LC - task id 7 - min elapsed time 18us (13us)
 SDB1_FAST_SHUTTER - task id 1  SDB1_LC - task id 5  SDB2_SBE - task id 8
 EDB_PC - task id 2  EDB - task id 6  SDB1_FI - task id 9 - min elasped time 7us (1,2us}
 SQZ_CTRL - task id 3    

The proposal is to use the EDB Acl task to manage the EDB_OMC  lock and to use the  unused elapsed time for the EBD_OMC algorithms .

To allow this a the task repartition on the 3 RT_CPUs has been reorganized as follows

RT_CPU2 RT_CPU1  RT_CPU0
 SDB1_OMC - task id 0  SDB1_Quadrants - task id 4  SDB2_LC - task id 7 - min elapsed time 0us (13us)
 SDB1_FAST_SHUTTER - task id 1  SDB1_LC - task id 5  SDB2_SBE - task id 8
 EDB_PC - task id 2  SDB_FI - task id 6  EDB - task id 9 - min elasped time 13us (3us}
SQZ_CTRL - task id 3    

DAQ Boxes configuration: EDB_DBOX_DetLab and EDB_DBOX_DetLab2

  • EDB_DBOX_DetLab (DBOX_SN51)
    • EDB_ADC (ADC2378_SN06) : add 3 new ADC channels at channel Id 2,3,4 for respectively EDB_OMC1_PZT_out_raw (100KHz), EDB_OMC1_Peltier_out_raw(100KHz) and EDB_OMC1_Temp_out_raw(10KHz ) and sent to the SDB_EDB_rtpc
    • EDB_MezzPD (SERVICE_SN10):  add  2 new photodiodes EDB_B1t_PD1 on channel 0-1 acquired at 100KHz and EDB_B1s_PD1 on channel 4-5 acquired at 100KHz . To receive correctly these new channels, the EDB_B1s1 signals read on channel 2-3 are now acquired at 10KHz instead of 100KHz previously
  • EDB_DBOX_DetLab2 (DBOX_SN113)
    • EDB_DEMOD_B1s_B1t (DEMOD_SN43) add the 2 new photodiodes EBD_B1t _PD1 on channel id 0 and EDB_B1s_PD1 on channel id 1. These 2 RF photodiodes are then demodulated at F6MHz, 2xF6MHz, F56MHz and 2xF56MHz. To run
      • the sample and monitoring channels are sent to the SDB_EDB_rtpc(rtpc1). 
      • the data channels are sent to the SDB2_rtpc(rtpc13)

New channels readout

  • The EDB_B1s_PD and EDB_B1t_PD RF channels are managed by the SDB2_Readout server where the last  part  of the digital demodulation has been setup.
  • The EDB_B1s_PD and EDB_B1t_PD RF  Audio and DC channels  are managed by the EDB server, as well   the EDB_OMC1 part
    • the code related to the EDB_OMC1 part remains to be added in the EDB server

Note that for these 2 rtpcs, SDB_EDB_rtpc and SDB2_rtpc, we are close to the maximum data flow that can be handed over 1 single Tolm link : adding a second Tolm link  between the MxDx in the DAQ room and the ones in the CEB computing room could be helpfull

 

Comments to this report:
mwas - 18:06 Thursday 03 April 2025 (66505) Print this report

The work has change the sign of the OMC error signal. To recover the locking  I have changed by pi the phase of B1_PD3 and B1 demodulation for the OMC length dither line. I have changed it using the external file, and also changed for the future the main file. After that change the OMC relocked at the first attempt.

gouaty, masserot, was - 18:29 Thursday 03 April 2025 (66504) Print this report

Related to the EDB server , the code for the EDB_OMC1 control has been added successfuly . It remains to test it

Cabling informations

  • EDB_DBOX_DetLab - DaqBox SN52
    • mezzanine 0: DAC1955-SN46
      • ch02 - OMC1_PZT_dac : PZT driving signal
      • ch03 - OMC1_Peltier_dac : Peltier driving signal
    • mezzanine 2: ADC2378-SN06
      • ch02 - EDB_OMC1_PZT_out : PZT command readout
      • ch03 - EDB_OMC1_Peltier_out : Peltier command readout
      • ch04 - EDB_OMC1_Temp_out : OMC1 Temperature sensor readout
    • mezzanie 3 - SERVICE-SN10
      • ch00-01 - EDB_B1t _PD1 : B1t_PD1 Audio/DC readout
      • ch04-05 - EDB_B1s_PD1 : B1s_PD1 Audio/DC readout
  • EDB_DBOX_DetLab2 - DaqBox SN113
    • mezzanine 3 - DEMOD-SN43
      • ch00 - EDB_B1t_PD : EBD_B1t RF readout
      • ch01 - EDB_B1s_PD : EDB_B1s RF  readout
masserot - 14:58 Friday 04 April 2025 (66515) Print this report

As the EDB_B1s1 channels sample frequency has been modified from 100KHz to 10KHz  the SDB1_OMC server has been restarted at 2025-04-03-07h50m00-UTC

Search Help
×

Warning

×